Please help me figure out this function.
Let f(x)=x-(2/x), Determine f(3) and f(a+b)
Thanks
Renee

Answer by Nate(3500)

f(x)=x-(2/x)
f(3)=3-(2/3)
f(3)=(7/3)
f(x)=x-(2/x)
f(a+b)=a+b-(2/(a+b))